# How AI Ideas Affect the Creativity, Diversity, and Evolution of Human Ideas: Evidence From a Large, Dynamic Experiment

### v1

@@ Appendix B AUT Prompts
We conducted a permutation test of whether the limited-length zero shot prompt (‘ZeroShotLimit’) yielded more similar word counts to the human responses than the original zero shot prompt (‘ZeroShot’) from (Stevenson et al., 2022). Note that permutation tests are non-parametric, so they are robust to (e.g.) violations of normality. The observed difference is computed as:
$T^{obs}=$
$|Avg(\text{ZeroShot})-Avg(\text{HumanResponses})|-$
$|Avg(\text{ZeroShotLimit})-Avg(\text{HumanResponses})|$
where $Avg$ is the average word count of a response and the human responses are from Organisciak et al. (2022).
Under the null hypothesis, we assume prompts do not differ in respective word counts so we can swap the labels of ‘ZeroShot’ and ‘ZeroShotLimit’. To simulate draws from this null distribution, we permuted the labels of the source column 10000 times, each time calculating a new value for the difference of differences defined in $T$. The p-value is the proportion of permutations where the null $T$ is greater than or equal to the observed difference, $T^{obs}$. As in Ojala and Garriga (2009), we apply a conservative adjustment and add $1$ to the numerator and denominator (this means we never get a p-value of 0).
$\text{p}=\frac{[\sum_{i=1}^{n}I(T^{null}>=T^{obs})]+1}{n+1}$
This p-value indicates the likelihood of observing a difference as extreme as our observed difference if the null hypothesis—that prompts did not differ in word counts—is true. We conclude the observed difference is significant at the $p<0.001$ level.
@@ Appendix E Human vs AI Ideas
We compared a sample of 1500 ideas from our modified Stevenson prompt in the prompt experiment and a random sample of 1500 ideas from the Organisciak Dataset for our 5 items. For each set, we used the model’s predicted originality scores. Originality ranges from 1-5. Overall, ChatGPT ideas had higher ($\beta=0.62$, $t(2994)=22.49$, 95% CI = $[0.56,0.67]$) originality. In the paper, we refer to this as a 12 p.p difference since $0.62/5=12\%$.
@@
sourcechatgpt ${}^{***}$
promptpants $*{}^{*}$
promptshoe ${}^{**}$
@@ I.1. Global
For 50 Monte Carlo runs with different seed values, we sampled 50 ideas for each {[condition], item} combination. For each 50-idea set, we computed various idea diversity measures. First, we calculated all pairwise SBERT distances. Next, we measured the mean, median, and max pairwise distance. We also computed the centroid of each 50-idea set and calculated the mean distance from the centroid. After calculating these metrics, we conducted two-tailed, paired permutation tests (10,000 iterations) to test if two conditions differed on these metrics. To conduct the paired permutation test, we randomly swapped the sign of the difference between pairs of values, equivalent to randomly swapping the condition labels of rows—simulating the null hypothesis that conditions do not differ. We then counted the proportion of null distribution iterations where one would observe a larger absolute difference in means than the observed difference. We added a 1 to the numerator and denominator, which is a common, conservative adjustment (Ojala and Garriga, 2009) and stops p-values from being 0. Because the test is paired (equivalent to swapping the condition label within each ‘row’), our permutation tests are controlling for both AUT items and Monte Carlo seeds, since each row shares these attributes. We controlled for multiple pairwise comparisons by applying a Holm-Bonferroni adjustment to p-values. As a non-parametric measure of effect size, we used Cliff’s Delta. This metric ranges from -1 to +1 where 0 indicates no difference between conditions, +1 indicates that all Monte Carlo runs for the first condition are larger than those for the second, and vice versa for -1. As with evolution, to reduce any effect of seeds, we consider ideas after the sixth trial (see Appendix I.3 for details.)
@@
0 HighExposureDisclosed-Control 1.29 0.0010 0.35
1 HighExposureUndisclosed-Control 0.97 0.0010 0.28
2 LowExposureDisclosed-Control 0.38 0.0066 0.10
3 LowExposureUndisclosed-Control -0.40 0.0748 -0.03
4 HighExposureDisclosed-HighExposureUndisclosed 0.32 0.0748 0.03
5 HighExposureDisclosed-LowExposureDisclosed 0.91 0.0010 0.27
6 HighExposureDisclosed-LowExposureUndisclosed 1.69 0.0010 0.30
7 HighExposureUndisclosed-LowExposureDisclosed 0.59 0.0010 0.22
8 HighExposureUndisclosed-LowExposureUndisclosed 1.37 0.0010 0.27
9 LowExposureDisclosed-LowExposureUndisclosed 0.78 0.0010 0.09
@@
0 HighExposureDisclosed-Control 1.18 0.0010 0.32
1 HighExposureUndisclosed-Control 0.87 0.0010 0.27
2 LowExposureDisclosed-Control 0.45 0.0024 0.12
3 LowExposureUndisclosed-Control -0.54 0.0086 -0.05
4 HighExposureDisclosed-HighExposureUndisclosed 0.31 0.0747 0.03
5 HighExposureDisclosed-LowExposureDisclosed 0.73 0.0010 0.24
6 HighExposureDisclosed-LowExposureUndisclosed 1.72 0.0010 0.30
7 HighExposureUndisclosed-LowExposureDisclosed 0.42 0.0060 0.19
8 HighExposureUndisclosed-LowExposureUndisclosed 1.41 0.0010 0.27
9 LowExposureDisclosed-LowExposureUndisclosed 0.99 0.0010 0.13
@@
0 HighExposureDisclosed-Control 1.42 0.0010 0.35
1 HighExposureUndisclosed-Control 1.10 0.0010 0.28
2 LowExposureDisclosed-Control 0.40 0.0072 0.10
3 LowExposureUndisclosed-Control -0.36 0.1032 -0.03
4 HighExposureDisclosed-HighExposureUndisclosed 0.32 0.1032 0.03
5 HighExposureDisclosed-LowExposureDisclosed 1.02 0.0010 0.27
6 HighExposureDisclosed-LowExposureUndisclosed 1.79 0.0010 0.30
7 HighExposureUndisclosed-LowExposureDisclosed 0.70 0.0010 0.22
8 HighExposureUndisclosed-LowExposureUndisclosed 1.46 0.0010 0.27
9 LowExposureDisclosed-LowExposureUndisclosed 0.76 0.0010 0.09
@@ I.2. Local
conditionLoExposure_Disclosed:creativity_human $*{}^{*}$ (0.018) ${}^{**}$ (0.017) ${}^{**}$ (0.028)
@@
conditionHiExposure_Undisclosed:ai_rel_create ${}^{**}$ (0.018) ${}^{**}$ (0.017) ${}^{**}$ (0.028)
@@
Bayesian Inf. Crit. 22,620.630 22,248.300 25,480.180
Note: $*{}^{*}$p$<$0.1; ${}^{**}$p$<$0.05; ${}^{***}$p$<$0.01
@@ I.3. Evolution
We took two additional steps to make sure our results were not driven by confounding factors. First, $Nobs$ controls for how many ideas are in the set that is being analyzed. Recall that we designed the experiment so that each (item, condition) combination was replicated exactly seven times in response chains of exactly 20 trials. However, there were some minor deviations (discussed in Appendix L) in response chains, resulting in some (item, condition, trial number) sets having fewer items than others. Hence, we control for the number of ideas in a set. Second, we only ran this analysis on data after the sixth trial in a response chain to rule out the effect of seeds on evolution. The logic here is that the condition with the most initial seeds (6) was the control condition. The experiment is designed to ‘shed’ all seeds after trial six since by that time there would have been six experiment responses, meaning the most recent six ideas in the control condition would now all be from the experiment, and hence no seeds present in the example sets. (Note that for all local analyses, we directly control for the number of seeds present in the example set as a fixed effect.)
@@
conditionLow ExposureDisclosed:trial_no 0.368 (0.233) $*{}^{*}$ (0.220) 0.196 (0.148)
@@
conditionHigh ExposureUndisclosed:trial_no ${}^{**}$ (0.239) ${}^{**}$ (0.225) ${}^{**}$ (0.151)
@@
conditionHigh ExposureDisclosed:trial_no ${}^{**}$ (0.239) ${}^{**}$ (0.225) ${}^{**}$ (0.151)
@@
conditionLoExposure_Undisclosed:creativity_human $*{}^{*}$ (0.029) 0.037 (0.023) 0.037 (0.023)
@@
conditionHiExposure_Disclosed:creativity_human ${}^{***}$ (0.029) ${}^{***}$ (0.023) ${}^{***}$ (0.023)
@@
conditionHiExposure_Undisclosed:creativity_human $*{}^{*}$ (0.029) 0.030 (0.023) $*{}^{*}$ (0.024)
@@
conditionHiExposure_Undisclosed:ai_rel_create ${}^{***}$ (0.028) ${}^{***}$ (0.022) ${}^{***}$ (0.023)
@@
conditionHiExposure_Undisclosed:interest_groupcreative $*{}^{*}$ (1.843) --2.329 (1.440) --2.344 (1.470)
@@
Initially, we did not implement any content moderation. But on July 8th, we received an influx of responses from Reddit. Several participants were trolls, providing repetitive profane responses. We then implemented a form of content moderation, flagging any idea that contained a word in a list of words banned by Google as of July 8, 2023131313https://github.com/coffee-and-fun/google-profanity-words and subsequently added two more words to the list. If an idea was flagged, it was written to our database but not shown to future participants. Some profane ideas were already shown to participants in between the time when the responses were submitted and we saw and implemented our solution. Also, the content moderation strategy was imperfect: 10 of 46 flagged ideas were false positives. As discussed earlier, condition was unrelated to the number of flagged ideas ($\chi^{2}$(4) = 6.06, p = 0.19) or total number of excluded ideas ($\chi^{2}$(4) = 3.87, p = 0.42).

### v2

@@ Appendix B AUT Prompts
We conducted a Monte Carlo experiment to confirm our modified zero-shot prompt resulted in responses with a similar word length to humans. Using parameters from Stevenson et al. (2022)’s experiments: For $n=1000$ trials, we fixed presence penalty and frequency penalty at 1, randomly chose a temperature (higher values lead to more randomness) in [0.65, 0.7, 0.75, 0.8], and randomly chose one of our 5 AUT items. For each trial, ChatGPT generated five ideas. The modified prompt resulted in responses with an average word length (M=4.44, SD = 1.34) much closer to human responses (M=4.56, SD = 4.97) than the original zero-shot prompt (M=25.38, SD=8.55). A permutation test further shows that this difference in word count was significant at $p<0.001$. We used the ideas generated by our modified zero-shot prompt as stimuli for the main experiment.
@@ Appendix E Human vs AI Ideas
We compared a sample of 1500 ideas from our modified Stevenson prompt in the prompt experiment and a random sample of 1500 ideas from the Organisciak Dataset for our 5 items. For each set, we used the model’s predicted originality scores. Originality ranges from 1-5. Overall, ChatGPT ideas had higher ($\beta=0.62$, $t(2994)=22.49$, 95% CI = $[0.56,0.67]$) originality.
@@
sourcechatgpt 0.618∗∗∗
promptpants --0.072
promptshoe 0.096∗∗
@@ I.1. Global
For 50 Monte Carlo runs with different seed values, we sampled 50 ideas for each {[condition], item} combination. For each 50-idea set, we computed various idea diversity measures. First, we calculated all pairwise SBERT distances. Next, we measured the mean and median pairwise distances. We also computed the centroid of each 50-idea set and calculated the mean distance from the centroid. After calculating these metrics, we conducted two-tailed, paired permutation tests (10,000 iterations) to test if two conditions differed on these metrics. To conduct the paired permutation test, we randomly swapped the sign of the difference between pairs of values, equivalent to randomly swapping the condition labels of rows—simulating the null hypothesis that conditions do not differ. We then counted the proportion of null distribution iterations where one would observe a larger absolute difference in means than the observed difference. We added a 1 to the numerator and denominator, which is a common, conservative adjustment (Ojala and Garriga, 2009) and stops p-values from being 0. Because the test is paired (equivalent to swapping the condition label within each ‘row’), our permutation tests are controlling for both AUT items and Monte Carlo seeds, since each row shares these attributes. We controlled for multiple pairwise comparisons by applying a Holm-Bonferroni adjustment to p-values. As a non-parametric measure of effect size, we used Cliff’s Delta. This metric ranges from -1 to +1 where 0 indicates no difference between conditions, +1 indicates that all Monte Carlo runs for the first condition are larger than those for the second, and vice versa for -1. As with evolution, to avoid the confounding effect of conditions differing in the number of seeds, we consider ideas after the sixth trial (see SI I.3 for a more detailed discussion). This is because the experiment is designed to ‘shed’ all seed examples after trial six.
@@
0 HighExposureDisclosed-Control 1.340000 0.001000 0.350000
1 HighExposureUndisclosed-Control 0.900000 0.001000 0.270000
2 LowExposureDisclosed-Control 0.340000 0.011100 0.090000
3 LowExposureUndisclosed-Control -0.460000 $0$0.020000
4 HighExposureDisclosed-HighExposureUndisclosed 0.440000 0.011100 0.050000
5 HighExposureDisclosed-LowExposureDisclosed 1.000000 0.001000 0.290000
6 HighExposureDisclosed-LowExposureUndisclosed 1.800000 0.001000 0.320000
7 HighExposureUndisclosed-LowExposureDisclosed 0.560000 0.001000 0.210000
8 HighExposureUndisclosed-LowExposureUndisclosed 1.360000 0.001000 0.280000
9 LowExposureDisclosed-LowExposureUndisclosed 0.800000 0.001000 0.090000
@@
0 HighExposureDisclosed-Control 1.210000 0.001000 0.310000
1 HighExposureUndisclosed-Control 0.810000 0.001000 0.260000
2 LowExposureDisclosed-Control 0.470000 0.001000 0.110000
3 LowExposureUndisclosed-Control -0.610000 $0$0.060000
4 HighExposureDisclosed-HighExposureUndisclosed 0.410000 0.032800 0.040000
5 HighExposureDisclosed-LowExposureDisclosed 0.740000 0.001000 0.230000
6 HighExposureDisclosed-LowExposureUndisclosed 1.820000 0.001000 0.320000
7 HighExposureUndisclosed-LowExposureDisclosed 0.330000 0.032800 0.180000
8 HighExposureUndisclosed-LowExposureUndisclosed 1.410000 0.001000 0.290000
9 LowExposureDisclosed-LowExposureUndisclosed 1.080000 0.001000 0.150000
@@
0 HighExposureDisclosed-Control 1.480000 0.001000 0.350000
1 HighExposureUndisclosed-Control 1.030000 0.001000 0.270000
2 LowExposureDisclosed-Control 0.360000 0.013800 0.090000
3 LowExposureUndisclosed-Control -0.430000 $0$0.020000
4 HighExposureDisclosed-HighExposureUndisclosed 0.450000 0.013800 0.050000
5 HighExposureDisclosed-LowExposureDisclosed 1.120000 0.001000 0.290000
6 HighExposureDisclosed-LowExposureUndisclosed 1.910000 0.001000 0.320000
7 HighExposureUndisclosed-LowExposureDisclosed 0.680000 0.001000 0.210000
8 HighExposureUndisclosed-LowExposureUndisclosed 1.460000 0.001000 0.280000
9 LowExposureDisclosed-LowExposureUndisclosed 0.790000 0.001000 0.090000
@@ I.2. Local
We found mixed evidence that belief in AI’s relative creativity moderates local idea diversity. Regression results showed a small but significant interaction effect between the [High Exposure, Undisclosed] condition and relative AI creativity ($\beta=0.038$, $t(3244)=2.149$, 95% CI = $[0.003,0.073]$, $p=0.03$). We probed this effect with estimated marginal means, predicting local idea diversity for the bottom and top decile of participants by perception of AI creativity. Top-decile participants had slightly higher local idea diversity than bottom-decile participants in the [High Exposure, Undisclosed] condition ($\Delta=2.62,d=0.34$) but although this difference was significant before multiple comparisons ($p=0.01$), it was not significant after adjusting for multiple comparisons, ($p=0.06$; see Appendix Table 17). Hence, we conclude there is mixed evidence for the role of belief in AI’s relative creativity as a moderator of local idea diversity.
@@
conditionLoExposure_Disclosed:creativity_human 0.034 (0.018) 0.040∗∗ (0.017) 0.070∗∗ (0.028)
@@
conditionHiExposure_Undisclosed:ai_rel_create 0.038∗∗ (0.018) 0.040∗∗ (0.017) 0.065∗∗ (0.028)
@@
Bayesian Inf. Crit. 22,620.630 22,248.290 25,480.180
Note: p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01
@@ I.3. Evolution
We took two additional steps to make sure our results were not driven by confounding factors. First, $Nobs$ controls for how many ideas are in the set that is being analyzed. Recall that we designed the experiment so that each (item, condition) combination was replicated exactly seven times in response chains of exactly 20 trials. However, there were some minor deviations (discussed in SI L) in response chains, resulting in some (item, condition, trial number) sets having fewer items than others. Hence, we control for the number of ideas in a set. Second, we only ran this analysis on data after the sixth trial in a response chain to rule out the effect of seeds on evolution. The logic here is that the condition with the most initial seeds (6) was the control condition. The experiment is designed to ‘shed’ all seeds after trial six since by that time there would have been six experiment responses, meaning the most recent six ideas in the control condition would now all be from the experiment, and hence no seeds present in the example sets. (Note that for all local analyses, we directly control for the number of seeds present in the example set as a fixed effect.)
@@
conditionLow ExposureDisclosed:trial_no 0.368 (0.233) 0.379 (0.220) 0.196 (0.148)
@@
conditionHigh ExposureUndisclosed:trial_no 0.525∗∗ (0.239) 0.461∗∗ (0.225) 0.335∗∗ (0.151)
@@
conditionHigh ExposureDisclosed:trial_no 0.566∗∗ (0.239) 0.561∗∗ (0.225) 0.378∗∗ (0.151)
@@
conditionLoExposure_Undisclosed:creativity_human 0.053 (0.029) 0.037 (0.023) 0.037 (0.023)
@@
conditionHiExposure_Disclosed:creativity_human 0.115∗∗∗ (0.029) 0.066∗∗∗ (0.023) 0.061∗∗∗ (0.023)
@@
conditionHiExposure_Undisclosed:creativity_human 0.050 (0.029) 0.030 (0.023) 0.039 (0.024)
@@
conditionHiExposure_Undisclosed:ai_rel_create --0.074∗∗∗ (0.028) --0.065∗∗∗ (0.022) --0.071∗∗∗ (0.023)
@@
conditionHiExposure_Undisclosed:interest_groupcreative --3.224 (1.843) --2.329 (1.440) --2.344 (1.470)
@@
Initially, we did not implement any content moderation. But on July 8th, we received an influx of responses from Reddit. Several participants were trolls, providing repetitive profane responses. We then implemented a form of content moderation, flagging any idea that contained a word in a list of words banned by Google as of July 8, 2023121212https://github.com/coffee-and-fun/google-profanity-words and subsequently added two more words to the list. If an idea was flagged, it was written to our database but not shown to future participants. Some profane ideas were already shown to participants in between the time when the responses were submitted and we saw and implemented our solution. Also, the content moderation strategy was imperfect: 10 of 46 flagged ideas were false positives. As discussed earlier, condition was unrelated to the number of flagged ideas ($\chi^{2}$(4) = 6.06, p = 0.19) or total number of excluded ideas ($\chi^{2}$(4) = 3.87, p = 0.42).